# Memorization, Emergence, and Explaining Reversal Failures: A Controlled Study of Relational Semantics in LLMs

### v1

@@ Abstract
Autoregressive LLMs perform well on relational tasks that require linking entities via relational words (e.g., father/son, friend), but it is unclear whether they learn the logical semantics of such relations (e.g., symmetry and inversion logic) and, if so, whether reversal-type failures arise from missing relational semantics or left-to-right order bias. We propose a controlled Knowledge Graph-based synthetic framework that generates text from symmetric/inverse triples, train GPT-style autoregressive models from scratch, and evaluate memorization, logical inference, and in-context generalization to unseen entities to address these questions. We find a sharp phase transition in which relational semantics emerge with sufficient logic-bearing supervision, even in shallow (2–3 layer) models, and that successful generalization aligns with stable intermediate-layer signals. Finally, order-matched forward/reverse tests and a diffusion baseline indicate that reversal failures are primarily driven by autoregressive order bias rather than deficient inversion semantics.
@@
Yihua Zhu1,3    Qianying Liu3    Jiaxin Wang1    Fei Cheng1    Chaoran Liu3 Akiko Aizawa2,3   Sadao Kurohashi1,3   Hidetoshi Shimodaira1,4 1Kyoto University   2University of Tokyo   3NII LLMC   4RIKEN {zhu.yihua.22h@st, wang.jiaxin.77y@st, feicheng@i, kuro@i, shimo@i}.kyoto-u.ac.jp {ying, cliu, aizawa}@nii.ac.jp
@@ 1 Introduction
While RQ1 concerns whether relational semantics can be learned at all, RQ2 focuses on how such knowledge—when present—is deployed at inference time. Recent work reports the reversal curse berglund2023reversal: a model trained on relational statements of the form “$A$ is the father of BB” often fails to answer the logically equivalent reversed query “$B$ is the son of AA.” This raises a fundamental ambiguity: are such failures caused by missing explicit inversion relational representations, or by order bias induced by left-to-right AR decoding? This leads to our second research question: RQ2: When a model has internalized relational semantics, are reversal-type failures primarily due to deficiencies in relational semantics, or to order bias in AR generation? Complementary, RQ1 and RQ2 intersect at a central issue: in LLMs, whether relation logic, represented by relational words, is in a form that supports robust, direction-agnostic inference.
Answering these questions using standard web-scale pretrained LLMs is w. Relational words are often polysemous and context-dependent, and uncontrolled pretraining data introduces the risk of contamination, making it difficult to distinguish genuine logical generalization from memorization allen2023physics.
@@
    Revisiting the reversal curse. We provide evidence that reversal-type failures primarily stem from order bias in left-to-right autoregressive decoding rather than missing relational semantics, and show that this issue is mitigated by bidirectional exposure during training. In contrast, diffusion-based LMs do not exhibit such behavior.
@@ 2 Methodology
This section defines our controlled experimental setup, which shown in Figure 2.
@@ 2.1 Controllable KG Synthetic Corpus
We construct a fully controllable synthetic corpus grounded in a KG schema. We show end-to-end text generation example in Table 3.
KG and Triple Generation
@@
Pre-train Corpus Construction
@@ 2.2 Training Setup
After pre-training, we perform SFT ouyang2022training on the SFT corpus (generated from train content only) to equip the model with question-answering capability. No experimental variables are introduced during SFT. More details about training setup see Appendix A.1.
@@ 2.3 Evaluation Setup
In-context learning logic QA
@@ 3 Result 1: Can auto-regressive LMs Memorize and Perform Basic QA?
4 Result 2: Can auto-regressive LMs learn relational word logical semantics, and when?
@@
For prompts in $\text{Q}_{\text{ICL-C}}$ (e.g., “A is the father of B. B is the son of ”), we extract, at each layer $l$, the next-token logit, softmax probability, and rank of the correct first token. We then compute layer-wise means over 1,800 prompts. (full setup are provided in the appendix A.4.2)
@@
AR LMs can learn logical semantics of relational word.
@@
We observe a phase transition phenomenon in $\text{Q}_{\mathrm{ICL}}$ once the number of inversion/symmetry training samples exceeds a threshold. This is consistent with prior reports of emergent capabilities with scale wei2022emergent, but we suggests that the effective trigger can be the scale of targeted logical evidence rather than model size alone. A plausible explanation is a strategy shift: below the threshold the model relies on local co-occurrence or weak memorization, whereas above it the logic signal is frequent and diverse enough to support a reusable rule mechanism, in line with grokking-style dynamics nanda2023progress. The opposite trends of $\text{Q}_{\mathrm{Logic}}$ (decreasing) and $\text{Q}_{\mathrm{ICL}}$ (increasing) with larger $N$ are consistent with $\text{Q}_{\mathrm{Logic}}$ being more affected by entity-memorization interference, while $\text{Q}_{\mathrm{ICL}}$ more directly reflects relational-rule induction.
@@
The Takeaway 4.3 from our layer-wise analysis matches the residual-stream/circuits view in which intermediate layers are the primary locus where features are written and composed into task mechanisms, while later layers mainly read out these mechanisms and align them to the output distribution elhage2021mathematical. Accordingly, when relational logic is truly learned, the corresponding mechanism can be formed early and then amplified consistently; when it is weak or not fully integrated, late layers may rely on unstable heuristics and the final readout can suppress the partial logic signal due to competition with other features or distributional alignment. Combined with mechanistic grokking accounts nanda2023progress, this suggests that failing models remain in a memorization/weak-mechanism regime, producing the observed “improve-then-drop” behavior near the output.
@@
Figure 7 shows how $Q_{\text{Logic}}$ and $Q_{\text{ICL}}^{\text{Comp}}$ change with training steps $E$ under fixed $N=20000$, $L=12$, and $K=10$. Panel(a) reveals a clear grokking transition: $\text{Q}_{\text{ICL-C}}^{\text{Inv}}$ and $\text{Q}_{\text{ICL-C}}^{\text{Sym}}$ abruptly jump to near-perfect accuracy around $E\approx 20$. Notably, this ICL Logic QA generalization, which is largely decoupled from direct memorization, emerges earlier than the improvement in memorization-dependent Logic QA. This ordering contrasts with the standard characterization of grokking as delayed generalization after memorizationwang2024grokked; power2022grokking; wu2025rote, and instead suggests that for relational semantics in our setting, generalization can precede memorization.
@@ 5.1 Task Setting
Nuance ($\text{Q}_{\text{logic}}$). If the evaluation content provides only ‘A is the father of B” ($A\!\rightarrow\!\texttt{father}\!\rightarrow\!B$), the inversion logic query $\text{Q}_{\text{Logic}}^{\text{Inv}}$ (‘Who is the son of A?” $\Rightarrow$ B) matches the training order ($\text{A}\!\rightarrow\!\cdot\!\rightarrow\!\text{B}$). Therefore, it is a Forward query with respect to information order, rather than a Reverse query like $\text{Q}_{\text{Mem}}^{\text{People}}$. The same logic applies to the symmetry query $\text{Q}_{\text{Logic}}^{\text{Sym}}$.
Task 1: One-directional training for relational attributes (people).
@@
Task 2: Bidirectional training for non-relational attributes (job).
For job facts, we include both forward and reverse surface templates in training (e.g., “chef is the job of A” and “A’s job is chef”) and evaluate $\text{Q}_{\text{Mem}}^{\text{Job}}$ using matched Forward/Reverse queries. Detail has been shown in Table 4.
@@ 5.3 Takeaway
Bidirectional Training substantially mitigates reversal failures.
With bidirectional training in Task 2 and diffusion model, Forward and Reverse accuracies on are nearly identical.
@@ 5.4 Discussion
Bidirectional Training Mitigates the Reversal Curse
We also observe a direct mitigation consistent with the order-based account. In Task 3, bidirectional training makes forward and reverse accuracies on $\text{Q}_{\text{non}}^{\text{job}}$ nearly identical, indicating that training exposure to both orders largely removes reversal failures. This complements prior mitigation efforts that adjust training signals lv2024analysis and is consistent with recent evidence that non-AR diffusion LMs are less sensitive to order reversal, further implicating left-to-right decoding bias as the root cause nie2025large.
@@ 6 Related Work
The reverse curse phenomenon was introduced by berglund2023reversal. Subsequent analyses attribute it to directional training signals induced by causal masking and propose training-time interventions that expose alternative conditioning patterns lv2024analysis; zhu2024towards. In parallel, LLaDAnie2025large suggests that departing from the standard AR paradigm can help mitigate this effect.
@@ 7 Conclusion
We train AR LMs from scratch with KG-based synthetic data to address two questions: (i) Can AR LMs learn relational word logical semantics, and when? (ii) What drives reversal-type failures: deficient relational semantics or order bias? In our results, we observe a sharp phase transition in which relational semantics emerge with sufficient logic-bearing supervision, even in shallow (2–3 layer) models, and successful generalization aligns with stable intermediate-layer signals. Moreover, order-matched forward/reverse tests indicate that reversal failures are primarily driven by AR order bias rather than deficient inversion semantics.
@@ Acknowledgements
This work was partially supported by JST SPRING JPMJSP2110 (YZ), JSPS KAKENHI 22H05106, 23H03355, and JST CREST JPMJCR21N3 (HS).
@@ A.1.2 Hyperparameter
We generate outputs with temperature (0.8) and (top_k=100).
@@ A.3 Data Generation Framework
We briefly described this in the Methodology 2, but the coverage was not sufficiently detailed or technical. In this section, we provide a more formal description of the KG-based corpus construction, along with detailed statistics for each dataset. In Table 3, we illustrate the end-to-end text generation pipeline with a concrete example.
@@ A.3.2 Training Corpus Construction
Here $D_{\alpha}$ denotes the training corpus and is generated by $\mathcal{P}_{\alpha}$ parameterized by $N=\texttt{num\_train\_samples}$, $K=\texttt{num\_template}$, and the fixed set of four sentence formats $\mathcal{F}=\{f^{(1)},f^{(2)},f^{(3)},f^{(4)}\}$(Detail examples has been shown in Table 4). The outer set $\{\cdot\}_{n=1}^{N}$ indexes the $N$ semantic samples, where the $n$-th sample is defined by a tuple of several KG triples $T^{(n)}_{i}$. For each semantic sample, we generate $K$ different paragraphs express same information, indexed by $\{\cdot\}_{k=1}^{K}$, each corresponding to a paragraph-level template choice. Within the $k$-th paragraph, each sentence $i\in\mathcal{I}_{\alpha}$ is produced by first drawing a format index $\tau^{(k)}_{i}$ independently and uniformly from $\{1,2,3,4\}$, then transforming triple $T^{(n)}_{i}$ with the selected format via $f^{(\tau^{(k)}_{i})}(T^{(n)}_{i})$. Finally, $\mathrm{Shuffle}(\cdot)$ randomly permutes the resulting several sentences to form the paragraph text, yielding $K$ shuffled paragraphs per semantic sample and $N$ samples in total for $D_{\alpha}$.

### v2

@@ Abstract
Autoregressive LLMs perform well on relational tasks that require linking entities via relational words (e.g., father/son, friend), but it is unclear whether they learn the logical semantics of such relations (e.g., symmetry and inversion logic) and, if so, whether reversal-type failures arise from missing relational semantics or left-to-right order bias. We propose a controlled Knowledge Graph-based synthetic framework that generates text from symmetric/inverse triples, trains GPT-style autoregressive models from scratch, and evaluates memorization, logical inference, and in-context generalization to unseen entities to address these questions. We find a sharp phase transition in which relational semantics emerge with sufficient logic-bearing supervision, even in shallow (2–3 layer) models, and that successful generalization aligns with stable intermediate-layer signals. Finally, order-matched forward/reverse tests and a diffusion baseline indicate that reversal failures are primarily driven by autoregressive order bias rather than deficient inversion semantics.
@@
Yihua Zhu1,3    Qianying Liu3    Jiaxin Wang1    Fei Cheng1    Chaoran Liu3 Akiko Aizawa2,3   Sadao Kurohashi1,3   Hidetoshi Shimodaira1,4 1Kyoto University   2University of Tokyo   3NII LLMC   4RIKEN {zhu.yihua.22h, wang.jiaxin.77y}@st.kyoto-u.ac.jp {feicheng, kuro, shimo}@i.kyoto-u.ac.jp {ying, cliu, aizawa}@nii.ac.jp
footnotetext: Corresponding author.Our code is available at https://github.com/YihuaZhu111/KG_synthetic_LLM_train.
@@ 1 Introduction
While RQ1 concerns whether relational semantics can be learned at all, RQ2 focuses on how such knowledge—when present—is deployed at inference time. Recent work reports the reversal curse berglund2023reversal: a model trained on relational statements of the form “$A$ is the father of BB” often fails to answer the logically equivalent reversed query “$B$ is the son of AA.” This raises a fundamental ambiguity: are such failures caused by missing explicit inversion relational representations, or by order bias induced by left-to-right AR decoding? This leads to our second research question: RQ2: When a model has internalized relational semantics, are reversal-type failures primarily due to deficiencies in relational semantics, or to order bias in AR generation? Taken together, RQ1 and RQ2 converge on a central question: in LLMs, is the logic of relations encoded by relational words represented in a way that supports robust, direction-agnostic inference?
Answering these questions using standard web-scale pretrained LLMs poses two key challenges. Relational words are often polysemous and context-dependent, and uncontrolled pretraining data introduces contamination risks, making it difficult to distinguish genuine logical generalization from memorization allen2023physics.
@@
    Revisiting the reversal curse. We provide evidence that reversal-type failures primarily stem from order bias in left-to-right autoregressive decoding rather than missing relational semantics, and show that this issue is mitigated by bidirectional data exposure during training. In contrast, diffusion-based LMs do not exhibit such behavior.
@@ 2 Methodology
This section defines our controlled experimental setup, as shown in Figure 2.
@@ 2.1 Controllable KG synthetic corpus
We construct a fully controllable synthetic corpus grounded in a KG schema. We show an end-to-end text generation example in Table 5.
KG and triple generation.
@@
Pre-train corpus construction.
@@ 2.2 Training setup
After pre-training, we perform SFT ouyang2022training on the SFT corpus (generated from train content only) to equip the model with question-answering capability. No experimental variables are introduced during SFT. More details about the training setup are provided in Appendix A.1.
@@ 2.3 Evaluation setup
In-context learning logic QA.
@@ 3 Result 1: Can auto-regressive LMs memorize and perform basic QA?
4 Result 2: Can autoregressive LMs learn relational word logical semantics, and when?
@@ 4.1 Task setting
For prompts in $\text{Q}_{\text{ICL-C}}$ (e.g., “A is the father of B. B is the son of ”), we extract, at each layer $l$, the next-token logit, softmax probability, and rank of the correct first token. We then compute layer-wise means over 1,800 prompts. (the full setup is provided in appendix A.4.2)
@@
AR LMs can learn the logical semantics of relational word.
@@
We observe a phase transition phenomenon in $\text{Q}_{\mathrm{ICL}}$ once the number of inversion/symmetry training samples exceeds a threshold. This is consistent with prior reports of emergent capabilities with scale wei2022emergent, but we suggest that the effective trigger can be the scale of targeted logical evidence rather than model size alone. A plausible explanation is a strategy shift: below the threshold the model relies on local co-occurrence or weak memorization, whereas above it the logic signal is frequent and diverse enough to support a reusable rule mechanism, in line with grokking-style dynamics nanda2023progress. The opposite trends of $\text{Q}_{\mathrm{Logic}}$ (decreasing) and $\text{Q}_{\mathrm{ICL}}$ (increasing) with larger $N$ are consistent with $\text{Q}_{\mathrm{Logic}}$ being more affected by entity-memorization interference, while $\text{Q}_{\mathrm{ICL}}$ more directly reflects relational-rule induction.
@@
The Takeaway 4.3 from our layer-wise analysis matches the residual-stream/circuits view in which intermediate layers are the primary locus where features are written and composed into task mechanisms, while later layers mainly read out these mechanisms and align them to the output distribution elhage2021mathematical. This interpretation is also related to logit lens and tuned lens analyses, which decode intermediate hidden states into output-space predictions to trace how model beliefs evolve across layers nostalgebraist2020logitlens; belrose2023eliciting. Accordingly, when relational logic is truly learned, the corresponding mechanism can be formed early and then amplified consistently; when it is weak or not fully integrated, late layers may rely on unstable heuristics and the final readout can suppress the partial logic signal due to competition with other features or distributional alignment. Combined with mechanistic grokking accounts nanda2023progress, this suggests that failing models remain in a memorization/weak-mechanism regime, producing the observed “improve-then-drop” behavior near the output.
@@
Figure 7 shows how $Q_{\text{Logic}}$ and $Q_{\text{ICL}}$ change with training steps $E$ under fixed $N=20000$, $L=12$, and $K=10$. Panel (a) reveals a clear grokking transition: $\text{Q}_{\text{ICL-C}}^{\text{Inv}}$ and $\text{Q}_{\text{ICL-C}}^{\text{Sym}}$ abruptly jump to near-perfect accuracy around $E\approx 20$. Notably, this ICL Logic QA generalization, which is largely decoupled from direct memorization, emerges earlier than the improvement in memorization-dependent Logic QA. This ordering contrasts with the standard characterization of grokking as delayed generalization after memorization wang2024grokked; power2022grokking; wu2025rote, and instead suggests that for relational semantics in our setting, generalization can precede memorization.
Effect of template diversity on generalization and memorization.
In the main experiments, we study the effect of $K$ only under the small-data setting (N=2, [^0] N=2{,}000). To test whether template diversity still matters when $N$ is already large enough to support strong generalization, we additionally fix N=10, [^0] N=10{,}000 and vary $K\in\{1,3,5,10\}$. As shown in Table 1, both $Q_{\text{ICL}}^{\text{Inv}}$ and $Q_{\text{ICL}}^{\text{Sym}}$ are low when $K$ is small, improve substantially at $K=5$, and exceed $95\%$ at $K=10$. This shows that strong relational generalization requires not only enough training samples, but also sufficient template diversity. Moreover, Figure 4 (b) shows that $Q_{\text{Logic}}$ is also weak when $K$ is small, suggesting that template diversity affects memorization as well as generalization.
Template $K$ 1 3 5 10
$Q_{\text{ICL-C}}^{\text{Inv}}$ (%) 13.50 30.22 66.61 95.00
$Q_{\text{ICL-C}}^{\text{Sym}}$ (%) 18.67 52.00 89.56 97.00
Table 1: Effect of template diversity $K$ on in-context generalization when the number of training samples is fixed at N=10, [^0] N=10{,}000.
Why does $\text{Q}_{\text{Logic}}$ decrease at large $N$?
As shown in Figure 1, $\text{Q}_{\text{Logic}}$ decreases at large $N$. A likely reason is that larger $N$ requires memorizing more entity-level facts, especially many distinct names, while model size is fixed. Figure 5(a) further shows that increasing the number of layers substantially improves $\text{Q}_{\text{Logic}}$ at large $N$, suggesting that the drop mainly comes from limited model capacity rather than failed relational rule learning.
@@ 5.1 Task setting
Nuance ($\text{Q}_{\text{Logic}}$). If the evaluation content provides only “A is the father of B” ($A\!\rightarrow\!\texttt{father}\!\rightarrow\!B$), the inversion logic query $\text{Q}_{\text{Logic}}^{\text{Inv}}$ (“Who is the son of A?” $\Rightarrow$ B) matches the training order ($\text{A}\!\rightarrow\!\cdot\!\rightarrow\!\text{B}$). Therefore, it is a Forward query with respect to information order, rather than a Reverse query like $\text{Q}_{\text{Mem}}^{\text{People}}$. The same logic applies to the symmetry query $\text{Q}_{\text{Logic}}^{\text{Sym}}$.
Task 1: One-directional data exposure for relational attributes (people).
@@
Task 2: Bidirectional data exposure for non-relational attributes (job).
For job facts, we include both forward and reverse surface templates in training (e.g., “chef is the job of A” and “A’s job is chef”) and evaluate $\text{Q}_{\text{Mem}}^{\text{Job}}$ using matched Forward/Reverse queries. Details are shown in Table 6.
@@ 5.3 Takeaway
Bidirectional data exposure substantially mitigates reversal failures.
When the same autoregressive model is trained with both forward and reverse surface forms, the Forward and Reverse accuracies become nearly identical, indicating that the forward–reverse gap is largely removed.
@@ 5.4 Discussion
Bidirectional data exposure mitigates the reversal curse
We also observe a direct mitigation consistent with the order-based account. In Task 2, bidirectional data exposure makes forward and reverse accuracies on $\text{Q}_{\text{non}}^{\text{job}}$ nearly identical, indicating that training exposure to both orders largely removes reversal failures. This complements prior mitigation efforts that adjust training signals lv2024analysis and is consistent with recent evidence that non-AR diffusion LMs are less sensitive to order reversal, further implicating left-to-right decoding bias as the root cause nie2025large.
6 Result 4: How general is the observed emergence phenomenon beyond the KG-synthetic from-scratch setting?
6.1 Extension to pretrained models
Task setting.
To test whether our findings extend beyond from-scratch training, we conduct an additional experiment on pretrained GPT-2 small (12 layers $\times$ 12 heads). Starting from the pretrained checkpoint, we continue training the model using only the KG-synthetic pre-training corpus, without QA-style SFT, and evaluate inversion and symmetry under the same ICL completion setting as in the main experiments. We vary the number of training samples $N\in\{0,10\mathrm{k},20\mathrm{k},30\mathrm{k},50\mathrm{k}\}$, where $N=0$ denotes the pretrained model without continue training.
NN 0 10,000 20,000 30,000 50,000
$\text{Q}_{\text{ICL-C}}^{\text{Inv}}$ 6.00 20.23 35.34 37.53 98.89
$\text{Q}_{\text{ICL-C}}^{\text{Sym}}$ 8.00 23.37 37.28 38.34 99.32
Table 2: Completion-based ICL accuracy $\text{Q}_{\text{ICL-C}}$ of pretrained GPT-2 small after continued training on the KG-synthetic corpus with different numbers of training samples NN.
Observations.
Without adaptation to the KG-synthetic corpus ($N=0$), pretrained GPT-2 performs near chance. Accuracy rises gradually with more $N$, remains below $40\%$ up to $N=30\mathrm{k}$, and then jumps to about $99\%$ at $N=50\mathrm{k}$.
Takeaway and discussion.
The pretrained model shows the same phase-transition-like behavior as the from-scratch setting, suggesting that this pattern is not specific to random initialization on synthetic data. Instead, even pretrained autoregressive models require sufficient rule-consistent supervision to internalize and generalize relational semantics.
6.2 Extension to more complex relations
To test whether the emergence behavior in Result 2 extends beyond symmetry and inversion, we study a more complex compositional relation: $A$ is father of $B$, $B$ is father of $C$ $\Rightarrow$ AA is grandfather of $C$. Using the same KG-synthetic setup, we evaluate unseen-entity generalization with $Q_{\text{ICL-C}}$ without SFT. When N=50, [^0] N=50{,}000, accuracy remains near $0\%$; when N=100, [^0] N=100{,}000, it rises sharply to $47.6\%$. Although this is lower than symmetry/inversion (about $95\%$) and requires a larger $N$, the qualitative pattern is similar: the model first fails to generalize the rule, then improves sharply once sufficient rule-consistent supervision is provided. This suggests that the emergence phenomenon is not specific to symmetry or inversion, but can generalize to more complex logical relational structures, albeit with higher data requirements and lower final accuracy.
@@ 7 Related work
The reverse curse phenomenon was introduced by berglund2023reversal. Subsequent analyses attribute it to directional training signals induced by causal masking and propose training-time interventions that expose alternative conditioning patterns lv2024analysis; zhu2024towards. Lin et al. (lin2024delving) further show that reversal generalization is highly sensitive to the structure of training documents, and interpret this effect through an inherent bias in fact recalling. In parallel, LLaDA nie2025large suggests that departing from the standard AR paradigm can help mitigate this effect.
@@ 8 Conclusion
We train AR LMs from scratch with KG-based synthetic data to address two questions: (i) Can AR LMs learn relational word logical semantics, and when? (ii) What drives reversal-type failures? In our results, we observe a sharp phase transition in which relational semantics emerge with sufficient logic-bearing supervision, even in shallow (2–3 layer) models. Moreover, order-matched forward/reverse tests indicate that reversal failures are primarily driven by AR order bias rather than deficient inversion semantics.
@@
ICL evaluation shortcut
Although our ICL setting tests generalization to unseen entities, the prompts involve only a small number of entities and explicitly provide a forward relational statement in context. As a result, some questions may be solvable via shallow matching or local entity correspondence, without requiring a fully abstract understanding of relational semantics. Therefore, our ICL results should be interpreted as evidence of rule-like generalization beyond memorization, but not as a definitive test of pure relational-semantic understanding.
Limited pretrained-model evaluation
Our extension to pretrained GPT-2 small remains limited in scope. Specifically, we evaluate only ICL completion accuracy, and do not examine forward–reverse gaps or reversal-type failures in the pretrained setting. Therefore, while these results suggest that phase-transition-like emergence is not specific to from-scratch training, they do not establish whether the order-bias findings in Section 5 generalize to pretrained autoregressive models.
@@ Acknowledgements
This work was partially supported by JST SPRING JPMJSP2110 (YZ), JSPS KAKENHI 22H05106, 23H03355, and JST CREST JPMJCR21N3 (HS). It was also supported by the “Development Acceleration Use” program of ABCI 3.0, provided by AIST and AIST Solutions. We sincerely thank all reviewers for their valuable comments, constructive suggestions, and strong support during the rebuttal stage.
@@ A.1.2 Hyperparameters
We generate outputs with temperature 0.8 and top-k = 100.
@@ A.3 Data generation framework
We briefly described this in section 2 (Methodology), but the coverage was not sufficiently detailed or technical. In this section, we provide a more formal description of the KG-based corpus construction, along with detailed statistics for each dataset. In Table 5, we illustrate the end-to-end text generation pipeline with a concrete example.
@@ A.3.2 Training corpus construction
Here $D_{\alpha}$ denotes the training corpus and is generated by $\mathcal{P}_{\alpha}$ parameterized by $N=\texttt{num\_train\_samples}$, $K=\texttt{num\_template}$, and the fixed set of four sentence formats $\mathcal{F}=\{f^{(1)},f^{(2)},f^{(3)},f^{(4)}\}$(Detail examples are shown in Table 6). The outer set $\{\cdot\}_{n=1}^{N}$ indexes the $N$ semantic samples, where the $n$-th sample is defined by a tuple of several KG triples $T^{(n)}_{i}$. For each semantic sample, we generate $K$ different paragraphs expressing same information, indexed by $\{\cdot\}_{k=1}^{K}$, each corresponding to a paragraph-level template choice. Within the $k$-th paragraph, each sentence $i\in\mathcal{I}_{\alpha}$ is produced by first drawing a format index $\tau^{(k)}_{i}$ independently and uniformly from $\{1,2,3,4\}$, then transforming triple $T^{(n)}_{i}$ with the selected format via $f^{(\tau^{(k)}_{i})}(T^{(n)}_{i})$. Finally, $\mathrm{Shuffle}(\cdot)$ randomly permutes the resulting several sentences to form the paragraph text, yielding $K$ shuffled paragraphs per semantic sample and $N$ samples in total for $D_{\alpha}$.